OpenStudy (anonymous):

Make it single logarithmic method 1/2[logbase3(x^2-y^2)-log(x+y)]

OpenStudy (p0sitr0n):

logbase3(x 2 - y 2) = log (x 2 -y 2) / log3 1/2 ( log (x-y)(x+y) / ( x+y) log3 ) = 1/2 ( log (x-y) / log3 )

OpenStudy (anonymous):

is that over log base 3? @P0sitr0n